How do I justify each step in the equation:
5(n + 4) + 9n = (5n + 20 = 9n
= 5n + (20 + 9n)
= 5n + (9n + 20)
= (5n + 9n) + 20
= (5 + 9)n + 20
= 14n + 2o

Answer by richard1234(7193)

Just use all the distributive, commutative laws out there:

5(n + 4) + 9n = (5n + 20) + 9n (distribute property)
= 5n + (20 + 9n) (associative property)
= 5n + (9n + 20) (commutative property)
= (5n + 9n) + 20 (associative property)
= (5 + 9)n + 20 (factoring out n)
= 14n + 20 (5 + 9 = 14)